OpenStudy (anonymous):

Write the equation of the line that passes through (2, 4) and (1, –3) in slope-intercept form. A.) y = 7x – 10 B.) y = 7x + 10 C.) y = 7x + 18 D.) y = 7x – 18

OpenStudy (anonymous):

First find the slope using \[m = \frac{y_1 - y_2}{x_1 - x_2}\]

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

You know what to do from here?

OpenStudy (anonymous):

after i find the slope, no

OpenStudy (anonymous):

AAlright. After you get the slope, tell me what you get and I'll show you the next step.

OpenStudy (anonymous):

7

OpenStudy (anonymous):

Slope is right...

OpenStudy (anonymous):

y - y1 = m(x - x1) y - 4 = 7(x - 2) y - 4 = 7x - 14 add four to both sides y = 7x - 18

OpenStudy (anonymous):

So close! \[y = 7x - 10\]

OpenStudy (anonymous):

oh, i should have subtracted the 4.

OpenStudy (anonymous):

No. You're actually adding the 4. y - 4 = 7x - 14 y - 4 + 4 = 7x - 14 + 4 y = 7x + 4 - 14 y = 7x - 10

OpenStudy (anonymous):

Since 14 is negative, when you subtract, you go farther negative which is -18. When you add, you cancel some of the negative and it becomes -10.

OpenStudy (anonymous):

ok! i understand now.

OpenStudy (anonymous):

thank you :D

OpenStudy (anonymous):

np :)